OpenStudy (anonymous):

There is a rather simple equation in my textbook in Physics but I'm missing a minus somewhere. Some help will be appreciated.

OpenStudy (anonymous):

Here it is: \[-i \frac{ d }{ dx }f(x) = kf(x)\]

OpenStudy (anonymous):

So, I separate the variables to: \[\frac{ df(x) }{ f(x) }=-ik d(x)\] Which becomes \[e^{-ikx}+C\] However, they don't give a minus in the exponent.

ganeshie8 (ganeshie8):

\(\large \frac{ df(x) }{ f(x) }=-\frac{k}{i} d(x) \)

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

1/i = -i so negative cancels out

OpenStudy (primeralph):

You should recognize that 1/i = -i.

OpenStudy (anonymous):

aah, right

OpenStudy (anonymous):

yeah, I made a mistake and put i in the numerator

OpenStudy (anonymous):

thank you both!

OpenStudy (primeralph):

Ganeshieeeeeeeee!

ganeshie8 (ganeshie8):

you trolling @beccaboo333 :)

ganeshie8 (ganeshie8):

u wlc @suzanka !

OpenStudy (beccaboo333):

>_> ~leaves~ sorry guys